OpenStudy (anonymous):

What is the equation of the ellipse with vertices at (-25, 0), (25, 0) and co-vertices (0, -15), (0, 15)? Answers: x2/625+ y2/225 = 1 x2/225 +y2/625 = 1 x2/225 - y2/625 = 1 x2/625 - y2/225 = 1

OpenStudy (anonymous):

\[\frac{x^2}{25^2} + \frac{y^2}{15^2} = \frac{x^2}{625} + \frac{y^2}{225} = 1\]